40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Feb 22 18:47:51 2022, Ian MacMillan, Update, SUS, ETMY SUS Electronics Replacement ALL_TF_Graph.pdf20220222_SUSElectronicsReplacement.7z
    Reply  Fri Feb 25 14:00:08 2022, Ian MacMillan, Update, SUS, ETMY SUS Electronics Replacement IMG_0853.jpgIMG_0852.jpg
       Reply  Fri Feb 25 14:48:53 2022, Ian MacMillan, Update, SUS, ETMY SUS Electronics Replacement Before.jpgAfter.jpg
          Reply  Sat Feb 26 23:48:14 2022, Koji, Update, SUS, ETMY SUS Electronics Replacement PXL_20220226_023111179_2.jpg
             Reply  Mon Feb 28 15:51:07 2022, Ian MacMillan, Update, SUS, ETMY 1Y4 Electronics Replacement IMG_0865.jpgIMG_0867.jpgIMG_0868.jpgIMG_0866.jpg
                Reply  Thu Mar 3 04:11:36 2022, Koji, Update, SUS, ETMY 1Y4 Electronics Replacement PXL_20220303_082726693.jpgPXL_20220303_082752494.jpgPXL_20220303_082744464.jpg
                   Reply  Thu Mar 3 04:24:23 2022, Koji, Update, SUS, ETMY 1Y4 Electronics Replacement PXL_20220303_094421604.jpgPXL_20220303_094435127.jpg
                      Reply  Thu Mar 3 17:21:11 2022, Ian MacMillan, Update, SUS, ETMY 1Y4 Electronics Replacement 
                         Reply  Sat Mar 5 02:03:46 2022, Koji, Update, SUS, ETMY 1Y4 Electronics Replacement 6x
                            Reply  Mon Mar 21 14:22:52 2022, Anchal, Update, SUS, ETMY Alignmnet offsets needed to be changed 
       Reply  Tue Mar 1 19:26:24 2022, Koji, Update, SUS, ETMY SUS Electronics Replacement PXL_20220302_025723651.jpgPXL_20220302_030102033.MP.jpg
    Reply  Mon Mar 7 16:44:15 2022, Ian MacMillan, Update, SUS, ETMY SUS Electronics Replacement 6x
       Reply  Fri Mar 11 16:43:03 2022, Ian MacMillan, Update, SUS, ETMY SUS Electronics Replacement 
          Reply  Fri Mar 25 11:39:14 2022, Anchal, Update, SUS, ETMY SUS Electronics Replacement - Questions 
             Reply  Mon Mar 28 16:25:34 2022, Ian MacMillan, Update, SUS, ETMY SUS Electronics Replacement - Questions 
    Reply  Thu Mar 17 11:40:41 2022, Paco, Update, SUS, ETMY green PZT HV supply 
Message ID: 16703     Entry time: Sat Mar 5 02:03:46 2022     In reply to: 16699     Reply to this: 16738
Author: Koji 
Type: Update 
Category: SUS 
Subject: ETMY 1Y4 Electronics Replacement 

Oplev saga

Summary

- The new coil driver had not enough alignment range to bring the oplev beam back to the QPD center
- The coil driver output R was reduced from 1.2k to 1.2k//100 = 92.3 +/- 0.4 Ohm
- Now the oplev spot could be moved to the center of the QPD

- The damping gains (POS/PIT/YAW) and the oplev gains were reduced by a factor of 1/10.
- The damping and the oplev servos work now. Fine gain tuning is necessary.

To Do:
- DC value / TF measurements
- Adjust damping gains
- RFM issue
- Connection check
- Cable labeling


== Alignment Range ==

- Since c1auxey was removed, we no longer have C1:SUS-ETMY_PIT_COMM and C1:SUS-ETMY_YAW_COMM. At this moment, all the alignment is taken with the offset input from the fast real-time system via C1:SUS-ETMY_PIT_OFFSET and C1:SUS-ETMY_YAW_OFFSET.

- The oplev spot could not be moved on the center of the QPD without exceeding the DAC output range (~+ or -32000) for the coils. (Attachment 1)

- This is because the old system had a slow but large current range (Rout = 100) and a small current range for the fast control. Until we commission the new HV BIAS Driver, we have to deal with the large DC current with the HAM-A coil driver.

== Modification to the output resistances ==

The following units and the channels were modified. Each channel had a differential current driver and two output resistances of 1.2K. 100Ohm (OHMITE 43F100, 3W) wire wound resistors were added to them in parallel, making the resulting output R of ~92Ohm.

- ETMY HAM A Coil Driver 1: S2100622 (Attachments 2/3) CH1/2/3
- ETMY HAM A Coil Driver 2: S2100621 (Attachments 4/5) CH3

- This modification allowed me to align the oplev spot to the center of the QPD. C1:SUS-ETMY_PIT_OFFSET and C1:SUS-ETMY_YAW_OFFSE are +2725 (8%FS) and -2341 (7%FS), respectively.
- The previous alignment slider values were -0.9392 and 0.7615 (out of 10). These are the reasonable numbers, considering the change of the Rout from 100 to 92Ohm, and the sign flip.
(By the way, autoBurt files for c1auxex and c1auxey were not properly configured and the history of C1:SUS-ETM*_*_COMM was not recorded.)

== Damping Servos ==

- Now, the POS/PIT/YAW servos experience ~x10 gains. So temporarily these gains were reduced (POS 20->2, PIT 6->0.6, YAW 4->0.4) and the loops are stable when engaged.
- Also the gains of the OPLEV servos were reduced from -4.5 to -0.45. The loops are stable when engaged.

== Snapshot of the working condition ==

Attachent 6 shows the screenshot for the snapshot of the working condition.


To Do

- The damping servos were tested without proper PD whitening compensation.
  -> It turned out this is not necessary as our modified PD whitening has the pole and zero at the same freqs as before.

- Compare the DC values of the OSEM outputs and compensate for the gain increase by the "cts2um" filter.

- The end RTS suffers from the RFM issue. There is no data transmitted from the vertex to the end. I suspect we need to restart the c1rfm process. But this will likely suspend all the vertex real-time machines. Careful execution is necessary.

- c1iscey has all the necessary analog connections. But they are not tested. When we lock the green/IR cavity, we'll need them.

- The cable labeling is only half done.

Attachment 1: oplev_spot.jpg  1.523 MB  Uploaded Mon Mar 7 00:24:52 2022  | Hide | Hide all
oplev_spot.jpg
Attachment 2: PXL_20220305_090003837.jpg  787 kB  Uploaded Mon Mar 7 00:42:11 2022  | Hide | Hide all
PXL_20220305_090003837.jpg
Attachment 3: PXL_20220305_090023436.jpg  2.266 MB  Uploaded Mon Mar 7 00:42:46 2022  | Hide | Hide all
PXL_20220305_090023436.jpg
Attachment 4: PXL_20220305_091232290.jpg  1.055 MB  Uploaded Mon Mar 7 00:42:57 2022  | Hide | Hide all
PXL_20220305_091232290.jpg
Attachment 5: PXL_20220305_091306604.MP.jpg  1.756 MB  Uploaded Mon Mar 7 00:43:12 2022  | Hide | Hide all
PXL_20220305_091306604.MP.jpg
Attachment 6: Screenshot_2022-03-05_01-37-26.png  227 kB  Uploaded Mon Mar 7 16:50:03 2022  | Hide | Hide all
Screenshot_2022-03-05_01-37-26.png
ELOG V3.1.3-